# Generative AI-Driven Human Digital Twin in IoT-Healthcare: A Comprehensive Survey

### v1

@@ IV-C Personalized Rehabilitation
During this personalized rehabilitation training, machine learning-based approaches are usually utilized to enhance the monitoring and evaluation process in the VT, providing objective and personalized insights. This method relies on substantial high-quality data to obtain a robust and accurate machine learning model. This data should encompass a diverse range of exercises and rehabilitation contexts, as well as holistic patient profiles. However, acquiring such data poses significant obstacles, such as data availability and privacy concerns, especially in individual-level. To this end, GAI can be used to generate synthetic data in the context of personalized rehabilitation training for enhancing the performance of the virtual “coach” role of VT, as shown in Fig. 13. For instance, Mennella et al. in [142] proposed a diffusion model-based approach, pose-guided condition diffusion model, to generate synthetic data that mimicked realistic-looking human movements in a rehabilitation context. Specifically, the data generation framework used the latent diffusion model in combination with ControlNet [143], and was trained on a pre-labeled dataset of 6 rehabilitation exercises. The total 22,352 synthetic images can accurately capture the spatial consistency of human joint relationships for predefined exercise movements. Additionally, for enhancing the machine learning-based post-stroke rehabilitation assessment, Boukhennoufa et al. in [144] proposed a GAN-based approach to generate synthetic data for amplifying two post-stroke rehabilitation datasets. Specifically, the proposed approach incorporated a Siamese network and an additional discriminator to address original GAN mode collapse issue. Experimental results showed that the proposed approach could generate more diverse and realistic data, and improving classification models’ accuracy in post-stroke rehabilitation assessment.
@@ V Future Research Directions
V-A Deriving Energy-Efficient Scheme for GAI-driven HDT
@@ V-B Designing Human-centric Evaluation Metrics
V-C Accelerating Real-time Responsiveness of GAI-driven HDT
@@ VI Conclusions
In this survey, we have shed light on implementing GAI-driven HDT in IoT-healthcare. Specifically, we have introduced IoT-healthcare, and envisioned the potential of utilizing GAI-driven HDT. Then, we have reviewed the fundamentals of HDT and GAI and illustrated the framework of GAI-driven HDT. Furthermore, we have delved into the implementation of GAI-driven HDT, including GAI enabled data acquisition, communication, data management, digital modeling, and data analysis. Based on this, we presented the IoT-healthcare applications of GAI-driven HDT, including personalized health monitoring and diagnosis, personalized prescription, and personalized rehabilitation. Finally, we have outlined some future research directions.

### v2

@@ IV-C Personalized Rehabilitation
During this personalized rehabilitation training, machine learning-based approaches are usually utilized to enhance the monitoring and evaluation process in the VT, providing objective and personalized insights. This method relies on substantial high-quality data to obtain a robust and accurate machine learning model. This data should encompass a diverse range of exercises and rehabilitation contexts, as well as holistic patient profiles. However, acquiring such data poses significant obstacles, such as data availability and privacy concerns, especially in individual-level. To this end, GAI can be used to generate synthetic data in the context of personalized rehabilitation training for enhancing the performance of the virtual “coach” role of VT, as shown in Fig. 13. For instance, Mennella et al. in [168] proposed a diffusion model-based approach, pose-guided condition diffusion model, to generate synthetic data that mimicked realistic-looking human movements in a rehabilitation context. Specifically, the data generation framework used the latent diffusion model in combination with ControlNet [169], and was trained on a pre-labeled dataset of 6 rehabilitation exercises. To evaluate the performance of the proposed approach, authors adopted qualitative and quantitative evaluation methods. As a rehabilitation specialist with extensive expertise in biomechanical and motion analysis, the first author of this paper undertook the assessment of the generated images for qualitative evaluation. As for quantitative evaluation, authors adopted the Inception score (IS), the Frechet Inception distance (FID), the Structural similarity index (SSMI), and the detection score for person-class (DS-person). These metrics provided a comprehensive evaluation of the fidelity, quality, and visual resemblance of the generated images to the ground truth images. Experiment results showed that the synthetic datasets achieved high-quality performance, i.e., IS = 162.94, FID = 299.06, SSMI = 0.545, DS-person = 0.789.
@@ V Future Research Directions
V-A Deriving Energy-Efficient Scheme
@@ V-B Designing Human-centric Evaluation Metrics
V-C Accelerating Real-time Responsiveness
@@
V-D Seamless Integration of GAI-driven HDT with Existing Healthcare IT Ecosystems
The integration of GAI-driven HDT into existing healthcare IT ecosystems is a fundamental requirement for the successful practical implementation. In this process, numerous many interoperability issues, such as efficient data exchange and effective data integration between these two types of systems. Standardized application programming interfaces (APIs) and web services is a practical solution to these challenges. In this regard, one significant integration issue to overcome is the existence of various data structures and standards within these two types of systems. To confront this complexity, deploying uniform data formats, such as health level seven (HL7) and fast healthcare interoperability resources (FHIR), is advisable. These formats are widely recognizable within global healthcare IT infrastructures, and their use in GAI-driven HDT can aid it in communicating effectively with existing healthcare IT infrastructures, allowing a streamlined process of data exchange and integration. However, the adoption of HL7 and FHIR in GAI-driven HDT would arise many challenges. For example, both them are complex data standards, with a broad range of elements and structures, which will pose a challenge for the GAI-driven HDT in terms of data processing and understanding. Additionally, mapping the data from the GAI-driven HDT to these formats is another challenging task due to varying data types and intricate data relationships. Tackling these integration and interoperability issues would provide a clearer roadmap for the practical implementation of GAI- driven HDT in IoT-healthcare.
@@ VI Conclusions
In this survey, we have shed light on implementing GAI-driven HDT in IoT-healthcare. Specifically, we have introduced IoT-healthcare, and envisioned the potential of utilizing GAI-driven HDT. Then, we have reviewed the fundamentals of HDT and GAI and illustrated the framework of GAI-driven HDT. After that, we have analyzed the implementation of GAI-driven HDT:
1) GAI can assist the data acquisition in HDT, by generating diverse and highly realistic synthetic data, including synthetic physiological, medical imaging and motion data.
2) GAI can offer efficient communication schemes for HDT, including GAI-aid semantic communication and GAI-aid cross-modal communication.
3) GAI can be used in data management of HDT by generating the missed data for data imputation, and clean data for data denoising. Additionally, GAI can produce synthesized yet authentic HDT data largely, preserving the original data distribution, whilst ensuring data security and privacy.
4) GAI can empower digital modeling in HDT, including digital modeling of human cells, tissues and organs.
5) GAI can facilitate data analysis in HDT, including classification, segmentation, anomaly detection and prediction.
Based on this, we have further presented the IoT-healthcare applications of GAI-driven HDT, such as personalized health monitoring and diagnosis, personalized prescription, and personalized rehabilitation. Finally, we have outlined some future research directions.
Acknowledgments
This work was supported by the State Key Laboratory of Massive Personalized Customization System and Technology under grant No. H&C-MPC-2023-04-01, and the National Research Foundation, Singapore, and Infocomm Media Development Authority under its Future Communications Research & Development Programme, Defence Science Organisation (DSO) National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP-2020-019 and FCP-ASTAR-TG-2022-003), and Singapore Ministry of Education (MOE) Tier 1 (RG87/22), and the National Natural Science Foundation of China (NSFC) under Grants No. 62102099 and No. U22A2054, and Postgraduate Research & Practice Innovation Program of Jiangsu Province No.KYCX24_0596.